### v1

# Neural Internal Model Control: Learning a Robust Control Policy via Predictive Error Feedback

Abstract
Accurate motion control in the face of disturbances within complex environments remains a major challenge in robotics. Classical model-based approaches often struggle with nonlinearities and unstructured disturbances, while reinforcement learning (RL)-based methods can be fragile when encountering unseen scenarios. In this paper, we propose a novel framework, Neural Internal Model Control (NeuralIMC), which integrates model-based control with RL-based control to enhance robustness. Our framework streamlines the predictive model by applying Newton-Euler equations for rigid-body dynamics, eliminating the need to capture complex high-dimensional nonlinearities. This internal model combines model-free RL algorithms with predictive error feedback. Such a design enables a closed-loop control structure to enhance the robustness and generalizability of the control system. We demonstrate the effectiveness of our framework on both quadrotors and quadrupedal robots, achieving superior performance compared to state-of-the-art methods. Furthermore, real-world deployment on a quadrotor with rope-suspended payloads highlights the framework’s robustness in sim-to-real transfer. Our code is released at https://github.com/thu-uav/NeuralIMC.
I Introduction
Achieving stable and precise motion in complex, disturbed environments is essential for robots, such as quadrotors maintaining stable flight in wind fields and legged robots carrying loads over diverse terrains. This challenge, known as the robust control or adaptive control problem, has been extensively studied [1]. Classical adaptive control methods adjust control parameters in real-time to compensate for dynamic environmental changes, often using mathematical models of the system and disturbances [2, 3, 4, 5], while robust control methods utilize more informative signals to enhance robustness, like Internal Model Control [6]. Despite their guaranteed stability, these methods rely on accurately modeling system dynamics, which is challenging for nonlinear systems with unstructured disturbances. In contrast, reinforcement learning (RL) methods have shown remarkable success in continuous control across various robot morphologies, including skill learning for armed robots [7, 8], high-speed quadrotor racing [9, 10], and adaptive locomotion for legged robots [11, 12, 13, 14, 15]. Large-scale training with randomized dynamics and environmental parameters enables RL-based control methods to generalize across different conditions and environments [16, 17], allowing robots to perform reliably even in previously unseen and highly varying scenarios.
Figure 1: Predictive errors can be caused by mismatched system modeling and/or external disturbances.
There has been a recent trend towards integrating model-based control with RL-based neural control to harness the strengths of both worlds [18, 19, 13, 20]. One representative framework is to use predicted future states or estimated disturbances as an additional input signal to obtain a more informative model-free RL policy. Long et al. [13] employ unsupervised contrastive representation learning to train a predictive encoder capable of estimating the next state in the latent space. Conversely, Huang et al. [20] use a handcrafted $\mathcal{L}_{1}$ adaptive law [21] to explicitly estimate applied force disturbances, resulting in a robust trajectory tracking controller that can handle unexpected wind fields. Although promising, integrating model-based and RL-based controls faces significant challenges. Predictive models often struggle with variability when training data cannot cover real-world dynamics, and handcrafted disturbance estimators require precise tuning, limiting their feasibility for complex systems and adaptability to dynamic changes. For establishing a more general method, we need to develop a structure that effectively combines these controls to fully leverage their advantages and address these limitations.
Figure 2: The closed-loop NeuralIMC framework using an explicit predictive model to calculate the next body state, applirope to both quadrotors and legged robots.
In this paper, we propose a general and flexible framework that integrates model-based and RL-based control. By simplifying the model to focus on the robot’s main body and treating its dynamics as a rigid transformation, we bypass the high-dimensional nonlinearity of complex robot systems and environmental interactions. This simplification enables us to use an explicit predictive model without requiring parameter tuning, establishing a foundation for robust closed-loop control. We compute the predictive error—defined as the discrepancy between plant and model responses—as a feedback signal to the policy, enhancing robustness and generalizability through informative feedback and extensive training. Our framework, shown in Fig. 2, can be easily integrated with other learning algorithms by incorporating this feedback loop. To demonstrate the effectiveness and versatility of our method, we apply it to two prevalent robot morphologies: quadrotors, which demand precise 3D control, and quadrupedal robots, which require adaptive locomotion to navigate complex terrain interactions. Empirical results indicate that our framework can surpass recent state-of-the-art control methods in both domains. Furthermore, we conduct a real-world deployment on a quadrotor with various rope-suspended payloads to illustrate its robustness in sim-to-real transfer.
II Related Work
II-A Classical control and disturbance estimation
Classical adaptive controllers require estimating system parameters in a closed-loop manner to adjust control actions in real-time. $\mathcal{L}_{1}$ adaptive control has been widely used to estimate force disturbances in quadrotors [22, 23]. Neural-Fly [18] incorporated pre-trained representations to enhance rapid online adaptation through deep learning, achieving stable flight in previously unseen wind conditions. Internal Model Control (IMC) [6, 3] is a general structure for robust control that uses a predefined model to predict system responses and estimate disturbances. However, applying IMC to more complex and nonlinear systems is challenging due to the difficulty of crafting reliable models for such systems. Long et al. [13] further proposed Hybrid Internal Model, in which they leverage a learning-based internal model with unsupervised contrastive learning. The learned model serves as a predictive encoder for estimating the system responses in the latent space. Despite their advantages, these methods either rely on explicit disturbance estimation or are heavily influenced by prediction accuracy, which can significantly affect performance if compromised.
II-B Learning-based adaptive control
With the rapid advancement of highly parallel simulators [17, 24, 25], reinforcement learning (RL) methods are increasingly demonstrating powerful capabilities to achieve robust control, including quadrotors [26, 20] and legged robots [11, 12, 13, 14, 15]. Unlike classical model-based control, these learning-based methods often use model-free RL algorithms like PPO [27] to train neural controllers in large-scale randomized environments. Koryakovskii et al. [28] proposed to incorporate an RL-baesd policy to generate residual actions to compensate nonlinear model predictive control in model-mismatch cases. Rapid Motor Adaptation (RMA) [12] is a pure RL-based methods, which predicts environmental parameters in the latent space using state-action histories through a teacher-student training procedure. This approach has been extended to quadrotors [26], bipedal robots [29], and humanoid robots [15]. A follow-up study found that a dual-history architecture, leveraging both short-term and long-term input/output history, achieves superior performance with a straightforward one-stage training process [14]. Conversely, DATT [20] incorporates an $\mathcal{L}_{1}$ adaptation law [21] to estimate forces applied to quadrotors directly, which are then used as policy inputs. Despite its effectiveness with precise parameter tuning, this method is sensitive to handcrafted settings and challenging to generalize to complex systems.
III Methods
In this section, we present a general framework that bridges model-based and RL-based control through a simplified predictive model, named NeuralIMC. We start by introducing the basics of Internal Model Control (IMC) and our extension to integrate it with RL-based neural control (Sec. III-A). By leveraging the model-free RL algorithm PPO [27], we reduce the complexity and accuracy requirements typically associated with classical model-based control. Next, we detail how rigid transformation is used to predict the robot’s body state, establishing a general approach to estimate disturbances and incorporate them into the control policy (Sec. III-B). Finally, we explain the application of this framework to two specific robot morphologies: quadrotors and quadrupedal robots (Sec. III-C).
III-A Internal Model Control
(a) Classical Feedback Control
(b) Internal Model Control
Figure 3: Feedback structures of classical feedback control and Internal Model Control (IMC).
Internal Model Control (IMC) is a powerful structure for robust control, offering significant advantages over classical feedback control, particularly in managing model uncertainties and external disturbances. Unlike classical feedback control, which relies on real-time error correction based on the difference between desired and actual outputs, IMC incorporates a predictive model of the process, allowing it to anticipate and compensate for disturbances more effectively. The structures of classical feedback control and IMC are illustrated in Fig. 3, where $P$ and $\tilde{P}$ represent the plant and internal models, respectively. $Q$ is the IMC controller, and $C$ is the classical feedback controller. The reference signal, disturbance, and system output are denoted by $r$, $d$, and $y$, respectively. In IMC, the predicted response $\tilde{y}$ is subtracted from the actual response $y$, and the difference, $\tilde{d}=\tilde{y}-y$, is used as the input signal to the IMC controller. Generally, $\tilde{d}$ is non-zero due to modeling errors and unknown disturbances not accounted for in the model. Consequently, IMC can provide more robust performance by continuously adjusting the control input according to the discrepancy between the predicted and actual responses.
While the design procedure for IMC is straightforward when the plant system is easy to model [3], it becomes significantly more complex and challenging for systems with nonlinear dynamics or when accurate models are difficult to obtain [30, 31]. This complexity arises because methods like linearization or approximation often fail to capture the true behavior of nonlinear systems, leading to performance degradation. Such a dependency on precise modeling limits its practical applications in robotic systems, such as quadrotors with cable-suspended payloads and legged robots on complex terrains, where achieving a robust and accurate system model is inherently challenging.
To address these challenges, our method combines the strengths of the IMC structure with large-scale RL training. A simplified predictive model in IMC provides a rough but stable estimate of the system’s response. Instead of relying solely on precise system modeling, our approach uses large-scale randomized simulations to learn a robust control policy that compensates for inaccuracies and simplified-out complexities. Through extensive training, the RL algorithm optimizes the control policy conditioned on predictive errors, enhancing robustness and performance in real-world scenarios where exact modeling is impractical. The predictive error reveals signals about external disturbances and internal characteristics of nonlinear dynamics, enabling more informative control decisions and effective policy learning.
III-B Design of a Simplified Predictive Model
The predictive model is crucial in our framework, providing the baseline needed to detect and compensate for system discrepancies and disturbances. Designing an effective predictive model involves two main challenges: accurately estimating the system response and ensuring the model’s predictions closely match actual system behavior. Accurate estimation is essential for the feedback loop to reflect the true robot state, while alignment between predictions and real responses is critical for reliable performance under varying conditions.
To streamline state estimation across diverse robots, we model only the rigid body state, excluding actuator- or limb-specific states. For robots equipped with monocular or binocular cameras and built-in IMUs, numerous visual(-inertial) odometry methods provide robust state estimations [32, 33, 34, 35], which can be used as the actual system response within the IMC structure. By treating the robot as a rigid body and using Euclidean transformation to estimate the next state, we significantly reduce the modeling complexity, making it more practical and efficient.
Here, we first present a general approach for formulating the predictive model and calculating the error, followed by specific adaptations for different robot morphologies in the next section. We assume that the model inputs include velocity or acceleration commands, and define the current body state as $\mathbf{x}_{k}=[\mathbf{p}_{k},\mathbf{q}_{k},\mathbf{v}_{k},\boldsymbol{\omega}_{k}]\in\mathbb{R}^{13}$, where $\mathbf{p}_{k}$ is the position, $\mathbf{q}_{k}$ is the quaternion (orientation), $\mathbf{v}_{k}$ is the velocity, and $\boldsymbol{\omega}_{k}$ is the angular velocity. The time interval between two steps is denoted as $\Delta t$.
The next position and velocity of the body can be computed via Newton’s law [36]:
$\tilde{\mathbf{p}}_{k+1}$ $=\mathbf{p}_{k}+\mathbf{v}_{k}\Delta t+\frac{1}{2}\mathbf{a}_{k}\Delta t^{2},$ (1)
$\tilde{\mathbf{v}}_{k+1}$ $=\mathbf{v}_{k}+\mathbf{a}_{k}\Delta t.$ (2)
For rotation, we adopt a discrete update rule while handling singularities. First, we compute the change in angle $\Delta\boldsymbol{\theta}$ with the angular velocity $\boldsymbol{\omega}_{k}$ and angular acceleration $\dot{\boldsymbol{\omega}}_{k}$:
$\Delta\boldsymbol{\theta}$ $=\boldsymbol{\omega}_{k}\Delta t+\frac{1}{2}\dot{\boldsymbol{\omega}}_{k}\Delta t^{2}.$ (3)
Next, we follow [37] to apply the exponential map to project the rotation from $\mathbb{R}^{3}$ to $SO(3)$, obtaining the corresponding quaternion $\mathbf{q}_{\Delta\theta}$:
$\mathbf{q}_{\Delta\theta}$ ={[$\cos\left(\frac{\|\Delta\boldsymbol{\theta}\|}{2}\right),\frac{\sin\left(\frac{\|\Delta\boldsymbol{\theta}\|}{2}\right)}{\|\Delta\boldsymbol{\theta}\|}\cdot\Delta\boldsymbol{\theta}\right],\|\Delta\boldsymbol{\theta}\|\leq\sqrt[4]{\epsilon_{\text{machine}}}\\ \left[\cos\left(\frac{\|\Delta\boldsymbol{\theta}\|}{2}\right),\left(\frac{1}{2}+\frac{\|\Delta\boldsymbol{\theta}\|^{2}}{48}\right)\cdot\Delta\boldsymbol{\theta}\right],\text{otherwise}\end{cases}$ (4)
Finally, we can predict the next quaternion and angular velocity as follows, where $\otimes$ denotes quaternion multiplication:
$\tilde{\mathbf{q}}_{k+1}$ $=\mathbf{q}_{k}\otimes\mathbf{q}_{\Delta\theta},$ (5)
$\tilde{\boldsymbol{\omega}}_{k+1}$ $=\boldsymbol{\omega}_{k}+\dot{\boldsymbol{\omega}}_{k}\Delta t.$ (6)
We have established an explicit method to predict the next body state, assuming the required motion can be represented by commands at the acceleration or velocity level. Then, given the predicted body state $\tilde{\mathbf{x}}_{k+1}=\left[\tilde{\mathbf{p}}_{k+1},\tilde{\mathbf{q}}_{k+1},\tilde{\mathbf{v}}_{k+1},\tilde{\boldsymbol{\omega}}_{k+1}\right]$ and the actual body state $\mathbf{x}_{k+1}$, we compute the predictive error $\tilde{\mathbf{d}}_{k+1}$, which is then fed into the policy to provide effective feedback signals.
$\tilde{\mathbf{d}}_{k+1}=\left[\Delta\tilde{\mathbf{p}}_{k+1},1-\cos{\tilde{\Theta}},\Delta\tilde{\mathbf{v}}_{k+1},\Delta\tilde{\boldsymbol{\omega}}_{k+1}\right]$ (7)
Here, $\Delta\tilde{\mathbf{p}}_{k+1}$, $\Delta\tilde{\mathbf{v}}_{k+1}$, and $\Delta\tilde{\boldsymbol{\omega}}_{k+1}$ represent the differences in position, velocity, and angular velocity, respectively. We use the cosine distance to quantify the difference between orientations, where $\tilde{\Theta}$ is the angle difference between the two quaternions.
III-C Specifications for Common Robots
In this section, we further explain how to derive velocity or acceleration commands for various robots, including quadrotors and quadrupedal robots. We demonstrate how these commands align with existing control methods, ensuring seamless integration into the policy learning process.
III-C1 Quadrotors
Traditional quadrotor controllers often use multi-layered cascaded structures [38], offering flexible options for integrating neural controllers. Recent research shows that using collective thrust and bodyrate (CTBR) as the action space for neural controllers provides significant advantages [39, 20, 9]. CTBR-based quadrotor controllers exhibit strong resilience to dynamics mismatches, effective domain transfer, and high agility and performance. This action space design is inherently compatible with our framework. We can directly use the bodyrate command from the policy to update the rotation, while the linear acceleration is computed using the collective thrust to translate the robot. Consequently, the predictive error, calculated from the previous policy output and the latest successive body states, can be seamlessly integrated into the current policy input without further modifications to the learning algorithm.
III-C2 Quadrupedal robots
Unlike quadrotors, legged robots, particularly quadrupedal and bipedal robots, have garnered significant attention for RL-based controllers due to their higher control dimensions and the complexities of modeling interactions with diverse terrains [8, 11, 12, 14, 19]. Typically, a neural controller for legged robots generates target actuator positions to achieve the desired body velocity, including forward velocity ($v_{x}$), lateral velocity ($v_{y}$), and yaw rate ($\omega_{z}$). To integrate our predictive model into the control loop, we use the reference velocity command as the model input instead of the high-dimensional actuator motions. By differentiating two consecutive reference velocity commands, we obtain the linear acceleration in x and y axes, and the angular acceleration for the yaw angle. In this way, we can apply the rigid transformation by setting other acceleration terms to zero, ensuring seamless integration into the existing policy structure. This method is compatible with other algorithms by simply adding a plug-and-play feedback loop to augment policy inputs.
IV Experimental Results
To evaluate our framework’s performance and versatility, we conduct experiments on quadrotors and quadrupedal robots, comparing against state-of-the-art methods in each domain. Then, we perform ablation studies on the feedback structure and noise tolerance of the simplified predictive model. Finally, we present sim-to-real deployment results on a quadrotor with two rope-suspended payloads, showcasing its effectiveness in enhancing system robustness.
TABLE I: Parameter ranges for dynamics randomization on quadrotor. $K_{p}^{0}=[37,37,11]$
. Parameter Name Train Range Eval Range Mass ($M$) [0.142, 0.950] [0.114, 1.140] Arm Length ($L$) [0.046, 0.200] [0.037, 0.240] Width-to-$L$ Ratio ($\alpha_{w}$) [1.0, 1.414] [0.852, 1.732] Height-to-$L$ Ratio ($\alpha_{h}$) [0.577, 1.0] [0.268, 1.414] Motor Force Constant ($K_{f}$) [1.15e-7, 7.64e-6] [9.16e-8, 9.17e-6] Max Thrust-to-Weight Ratio [2.0, 3.5] [1.6, 4.2] Motor Drag Constant ($\kappa$) [0.0041, 0.0168] [0.0033, 0.0201] Motor Time Constant ($\tau_{c}$) [0.3, 0.5] [0.24, 0.6] P Gains ($K_{p}$) [-0.3,0.3] ×$\times K_{p}^{0}$ [-0.5, 0.5] ×$\times K_{p}^{0}$
TABLE II: Simulation results for tracking errors (unit: m) with quadrotors, averaged over five random seeds. Bold values denote the best results. Blue cells indicate performance under random wind fields, with percentages in brackets representing the relative change in error compared to static conditions. T and S refer to the teacher and student models for RMA, respectively, while DATT with GT indicates the use of ground-truth disturbance inputs instead of $\mathcal{L}_{1}$ estimates.
Config Method Smooth Trajectory Infeasible Trajectory
Circle Chained Poly 5-Point Star Zigzag
Train MPPI+RL 0.087 0.157 (+80.46%) 0.056 0.118 (+110.71%) 0.103 0.175 (+69.90%) 0.105 0.164 (+56.19%)
PPO 0.102 0.169 (+65.69%) 0.065 0.117 (+80.00%) 0.097 0.172 (+77.32%) 0.095 0.154 (+62.11%)
RMA (T) 0.071 0.080 (+12.68%) 0.043 0.047 (+9.30%) 0.087 0.094 (+8.05%) 0.074 0.082 (+10.81%)
RMA (S) 0.084 0.110 (+30.95%) 0.048 0.064 (+33.33%) 0.116 0.157 (+35.34%) 0.112 0.142 (+26.79%)
DATT ($\mathcal{L}_{1}$) 0.084 0.128 (+52.38%) 0.051 0.076 (+49.02%) 0.089 0.163 (+83.15%) 0.077 0.138 (+79.22%)
DATT (GT) 0.084 0.093 (+10.71%) 0.052 0.055 (+5.77%) 0.089 0.098 (+10.11%) 0.077 0.084 (+9.09%)
Ours 0.072 0.083 (+15.28%) 0.032 0.038 (+18.75%) 0.065 0.073 (+12.31%) 0.057 0.064 (+12.28%)
Eval MPPI+RL 0.224 0.321 (+43.30%) 0.171 0.251 (+46.78%) 0.262 0.364 (+38.93%) 0.265 0.358 (+35.09%)
PPO 0.265 0.351 (+32.45%) 0.183 0.273 (+49.18%) 0.257 0.374 (+45.53%) 0.265 0.367 (+38.49%)
RMA (T) 0.117 0.147 (+25.64%) 0.072 0.097 (+34.72%) 0.151 0.192 (+27.15%) 0.145 0.187 (+28.97%)
RMA (S) 0.280 0.360 (+28.57%) 0.174 0.243 (+39.66%) 0.358 0.424 (+18.44%) 0.349 0.417 (+19.48%)
DATT ($\mathcal{L}_{1}$) 0.137 0.523 (+281.75%) 0.094 0.354 (+276.60%) 0.169 0.585 (+246.15%) 0.160 0.559 (+249.38%)
DATT (GT) 0.139 0.167 (+20.14%) 0.094 0.119 (+26.60%) 0.178 0.197 (+10.67%) 0.156 0.189 (+21.15%)
Ours 0.113 0.135 (+19.47%) 0.054 0.073 (+35.19%) 0.103 0.134 (+30.10%) 0.101 0.132 (+30.69%)
IV-A Experiment Settings
IV-A1 Quadrotors
[…]
IV-A2 Quadrupedal robots
[…]
IV-B Experiments on Quadrotors
We compare our method for quadrotors against four baselines: MPPI+RL [28], PPO [27], RMA [12], DATT [20], using two variants each for RMA and DATT to enable deeper analyses. All methods use the same PPO training parameters. For MPPI+RL, MPPI serves as a sampling-based nonlinear MPC method, with the RL policy compensating the MPPI outputs. We set the sampling number to 4096 and the rollout horizon to 20 in MPPI’s sampling procedure. PPO employs a three-layer MLP policy and a temporal convolutional network (TCN) to process future waypoints in the reference trajectory, which is inherited by all other methods. In addition, RMA adopts a teacher-student training framework, with RMA(T) as the teacher model and RMA(S) as the student. RMA(T) integrates dynamic parameters and ground-truth force disturbances, while RMA(S) uses a TCN to approximate the teacher’s latent state from state-action histories. In contrast, DATT utilizes an $\mathcal{L}_{1}$ disturbance estimator and tracking error inputs to improve trajectory-tracking robustness. To separate the $\mathcal{L}_{1}$ estimator’s effect, we also include a privileged variant, DATT(GT), which directly uses ground-truth force disturbances in place of the estimator.
As shown in Tab. II, our method consistently outperforms others across nearly all tasks, particularly under disturbances and out-of-distribution system dynamics. Among the baselines, DATT shows strong resilience to external disturbances, supported by its $\mathcal{L}_{1}$ estimator. However, its performance declines noticeably with unexpected dynamics in evaluation. Comparing DATT(GT) and DATT($\mathcal{L}_{1}$) reveals that the learned RL policy generalizes well when provided with ground-truth inputs, suggesting that the performance drop is primarily due to the handcrafted $\mathcal{L}_{1}$ estimator, which struggles to generalize across different systems. On the other hand. the teacher-student training approach in RMA results in the student policy underperforming compared to the teacher policy and shows instability under unseen systems, highlighting its unreliability for learning-based prediction.
In contrast, our method employs the proposed explicit model to establish a baseline prediction of the expected body transformation, effectively detecting discrepancies caused by internal and/or external sources. This informative feedback enhances the control policy’s responsiveness. Consequently, our method outperforms all baselines, including privileged variants like DATT(GT) and RMA(T), likely due to the informative predictive error feedback.
TABLE III: Simulation results on quadrupedal robots, averaged over five random seeds. The tracking error for linear velocity and angular velocity are computed as $\|v_{x,y}-v_{x,y}^{\text{target}}\|_{2}$ (unit: m/s) and $\|\omega_{\text{yaw}}-\omega_{\text{yaw}}^{\text{target}}\|_{2}$ (unit: rad/s), respectively. The best results for each case are highlighted in bold.
Disturbance Method Slopes Rough Slopes Stairs Discrete Obstacles
Linear Angular Linear Angular Linear Angular Linear Angular
None PPO 0.132 0.085 0.143 0.087 0.200 0.085 0.163 0.085
HIM 0.100 0.067 0.103 0.068 0.181 0.067 0.147 0.067
Ours (w/ HIM) 0.077 0.052 0.079 0.053 0.126 0.053 0.105 0.052
Ours 0.071 0.048 0.073 0.049 0.101 0.048 0.085 0.048
Impulse PPO 0.151 0.101 0.162 0.103 0.216 0.101 0.178 0.101
HIM 0.117 0.080 0.120 0.081 0.186 0.080 0.157 0.080
Ours (w/ HIM) 0.091 0.061 0.093 0.061 0.139 0.061 0.117 0.061
Ours 0.086 0.058 0.087 0.059 0.115 0.059 0.099 0.058
IV-C Experiments on Quadrupedal Robots
For quadrupedal robots, we compare our methods with PPO [27] and HIM [13]. PPO serves as a standard baseline in continuous control for quadrupeds, while HIM leverages unsupervised contrastive learning to train a predictive encoder for the next latent state and body velocity. Our method enhances the PPO policy by integrating predictive error feedback. Additionally, we evaluate a variant of our method, adding predictive error feedback to both the inputs of the HIM policy and its predictive encoder during training and evaluation.
As shown in Tab. III, our method demonstrates superior performance, achieving relative improvements over HIM ranging from approximately 26% to 44% across various test conditions. The combination of our framework with HIM also shows improvements over the original HIM. Interestingly, the predictive encoder in HIM shows a slight degradation compared to standard PPO when combined with our NeuralIMC structure. We hypothesize that this is because the predictive error provides sufficient signal for robust control, rendering the additional impact of predictive encoding negligible. This suggests that predictive error feedback alone is sufficiently informative for enhancing control robustness, simplifying the neural control pipeline, and facilitating integration with other learning algorithms.
IV-D Ablation Studies
In this section, we conduct two ablation studies to further explore the proposed framework. First, inspired by [14], we investigate whether integrating short- and long-horizon input/output histories into the feedback structure can enhance performance. Next, we introduce noise into the internal predictive model to test its noise tolerance and assess the impact of prediction accuracy on control performance. We perform all experiments on quadrotors tracking zigzag trajectories across three random seeds.
(a) Varying short-history length.
(b) Varying long-history length.
Figure 4: Ablation study on the feedback structure with different history lengths.
TABLE IV: Evaluation of the internal predictive model’s noise tolerance, tested on quadrotors tracking zigzag trajectories.
Noise $\sigma_{n}$ Train Eval
static windy static windy
0. 0.057 0.064 (+12.28%) 0.101 0.119 (+17.82%)
0.02 0.068 0.077 (+13.24%) 0.136 0.172 (+26.47%)
0.04 0.077 0.103 (+33.77%) 0.161 0.230 (+42.86%)
0.08 0.083 0.130 (+56.63%) 0.201 0.285 (+41.79%)
0.16 0.087 0.150 (+72.41%) 0.240 0.343 (+42.92%)
(a) Quadrotor system for deployment.
(b) Performance across different disturbances.
(c) Effect of the predictive error feedback.
Figure 5: Real-world experiments on a quadrotor with two different rope-suspending payloads.
IV-D1 History structure
We evaluate different feedback structures incorporating historical records. In short-history experiments, we concatenate several recent input/output pairs, including predictive error within each pair, into a single input vector for the policy. For long-history experiments, we retain the latest state-action pair while providing additional historical context. Following prior work [12, 14], we utilize a TCN model to process long-horizon history into a latent vector, which is then concatenated with the current input for the policy.
As shown in Fig. 4, increasing the length of historical inputs in the short- or long-horizon period does not improve performance compared to using the immediate response (i.e., using only the lated one-step short history). Although this differs from [14], it aligns with classical adaptive control methods such as $\mathcal{L}_{1}$ adaptive control [21] and robust control methods like internal model control [2, 3], where immediate error feedback suffices for improving motion accuracy and robustness. This simpler structure eases implementation and avoids cumulative error issues often encountered in real robot systems.
IV-D2 Noise tolerance
To assess the impact of the internal predictive model’s accuracy on control performance, we introduce varying levels of noise into the model’s predictions. Specifically, we add zero-mean Gaussian noise with different variances to the responses of our simplified internal model. We train and evaluate multiple policies using a range of noisy predictive models for quadrotor tracking on zigzag trajectories, with the results presented in Tab. IV.
The results indicate that our method maintains competitive performance under slight noise, though the accuracy of the predictive model becomes increasingly critical for larger noise levels, particularly in scenarios with external disturbances or out-of-distribution system dynamics. Across different training and evaluation ranges, performance declines as noise intensifies, with a similar trend observed under windy force disturbances. This observation highlights an advantage of our explicit predictive model over other learning-based approaches: it provides a stable baseline estimation, albeit approximate. Although modeling noise can affect feedback reliability, the rigid-body transformation adheres to fundamental motion principles and generalizes across robotic systems. While this model simplifies nonlinear complexities, the RL part in our framework compensates effectively within an informative closed-loop system.
IV-E Real-world Deployment
To further validate the effectiveness of our framework, we conducted a real-world experiment on a quadrotor, deploying our method to assess its robustness under various disturbances. The quadrotor, equipped with an Nvidia Jetson Xavier NX [42] and a PX4 flight control unit (FCU) [43], interfaces with a motion capture system for real-time pose tracking. The whole neural control framework runs on the embedded computer. The control policy generates CTBR commands, transmitted to the FCU via MAVROS [44] and converted to motor velocities. The quadrotor weighs approximately 1 kg, and we evaluated its tracking performance across five trials under three disturbance conditions: no disturbance, a 200g payload, and a 330g payload, each attached via a rope to increase instability and nonlinearity. The deployment system setup is shown in Fig. 5(a).
We present the tracking performance for each axis, along with the normalized error across all axes, in Fig. 5(b). The lines represent the average tracking errors under different conditions, with black intervals indicating variance. Our results show that the framework maintains relatively stable performance across various payloads. To illustrate the impact of predictive error feedback, we include a detailed example trajectory for the 330g payload as an ablation study in a real-world scenario (Fig. 5(c)). The z-axis trajectory demonstrates that, without predictive error feedback, the control policy struggles to maintain a stable flight height due to the unexpected heavy payload. In contrast, our method, which integrates predictive error feedback, achieves robust tracking despite the downward force from the additional payload.
V Conclusion
In this paper, we propose a novel framework, Neural Internal Model Control (NeuralIMC), that combines model-based control with RL-based neural control. By focusing on body dynamics and adopting the white-box rigid-body transformation, we bypass high-dimensional nonlinearity and employ a simplified explicit predictive model to form a closed-loop feedback structure. Integrating model-free RL algorithms allows the neural control policy to react according to the predictive error feedback. Our method demonstrates superior performance on quadrotors and quadrupedal robots, surpassing state-of-the-art control methods in both domains. Real-world deployment on a quadrotor with heavy payloads highlights the robustness of our approach in handling disturbances and adapting to varying dynamics. This work underscores the potential of our combined framework for enhanced robustness and adaptability in robotic systems and suggests a direction for more interpretable learning-based control systems in future research.
Limitations. While extensive simulation experiments highlight the advantages of the proposed neural control framework, real-world evaluation remains limited. We plan to test our framework on a wider range of robots, including quadrotors of varying sizes and legged robots on complex terrains. Additionally, our method can be further strengthened by incorporating more effective disturbance training [45].
Appendix A Specification Details
In this section, we describe the integration of the simplified predictive model discussed in Sec. III-B into the neural control pipeline of two widely used robotic systems: quadrotors and quadrupedal robots. We specifically detail the process of obtaining the necessary model inputs to update the predicted body states. For clarity, we omit the subscript $k$; all states—position $\mathbf{x}$, quaternion $\mathbf{q}$, velocity $\mathbf{v}$, and angular velocity $\boldsymbol{\omega}$—are from the previous time step, i.e., $k=t-1$. Our goal is to predict the expected body state at the current time step $t$.
A-A Quadrotors
For quadrotors, as demonstrated in Sec. III-C, we define the action space of the policy using mass-normalized collective thrust and body rates (CTBR). This is followed by a PD controller to manage the quadrotor’s flight dynamics.
We denote the collective thrust as $\mathbf{a}_{z}^{\mathcal{B}}$ and the body rates as $\boldsymbol{\omega}^{\mathcal{B}}$, both expressed in the quadrotor’s body frame. The acceleration in the world frame can then be computed as follows:
$\mathbf{a}=\mathcal{T}_{\mathbf{q}}(\mathbf{a}_{z}^{\mathcal{B}})+\mathbf{g}$ (8)
where $\mathcal{T}_{\mathbf{q}}$ represents the rotation of the z-axis acceleration from the body frame to the world frame, and $\mathbf{g}=[0,0,-9.81]^{T}$ is the free-fall acceleration. Then, we can apply Eqs. 1 and 2 to obtain the predicted position $\tilde{\mathbf{x}}$ and velocity $\tilde{\mathbf{v}}$.
For the rotation, we set $\dot{\boldsymbol{\omega}}$ to zero and directly replace $\boldsymbol{\omega}$ with $\boldsymbol{\omega}^{\mathcal{B}}$. Then, we can apply Eqs. 3, 4, 5 and 6 to obtain the predicted quaternion $\tilde{\mathbf{q}}$ and angular velocity $\tilde{\boldsymbol{\omega}}$.
A-B Quadrupedal Robots
For quadrupedal robots, we utilize reference velocity commands as the model inputs rather than the policy outputs. Given that the provided velocity commands do not cover all degrees of freedom, we set the unspecified parts to zero. We then use the command values as target quantities, calculating the required acceleration by taking the difference between the target and the actual state of the robot, in order to update the state.
Given the reference forward velocity $\hat{v}_{x}^{\mathcal{B}}$, lateral velocity $\hat{v}_{y}^{\mathcal{B}}$, and yaw rate $\hat{\omega}_{z}^{\mathcal{B}}$, all defined in the body frame, we can compute the desired acceleration as follows:
$\mathbf{a}^{\mathcal{B}}=\begin{bmatrix}(\hat{v}_{x}^{\mathcal{B}}-v_{x}^{\mathcal{B}})/{\Delta t}\\ (\hat{v}_{y}^{\mathcal{B}}-v_{y}^{\mathcal{B}})/{\Delta t}\\ 0\end{bmatrix}$ (9)
In this context, $v_{x}^{\mathcal{B}}$ and $v_{y}^{\mathcal{B}}$ denote the forward and lateral velocities of the body from the previous step, respectively, transformed into the body frame. The world-frame acceleration can then be obtained by applying the transformation $\mathcal{T}_{\mathbf{q}}$:
$\mathbf{a}=\mathcal{T}_{\mathbf{q}}(\mathbf{a}^{\mathcal{B}})$ (10)
Similarly, the angular acceleration can be computed as:
$\dot{\boldsymbol{\omega}}=\begin{bmatrix}0\\ 0\\ \hat{\omega}_{z}^{\mathcal{B}}-\omega_{z}^{\mathcal{B}}\end{bmatrix}$ (11)
Subsequently, the computed acceleration $\mathbf{a}$ and angular acceleration $\dot{\boldsymbol{\omega}}$ can be substituted into the relevant equations (Eqs. 1, 2, 3, 4, 5 and 6) to predict the next body state $\tilde{\mathbf{x}}$.
Appendix B Experiment Details
In this section, we provide a detailed explanation of how we set up the simulation tasks for training and evaluating various controllers for both quadrotors and quadrupedal robots.
B-A Quadrotors
B-A1 Quadrotor Dynamics
We employ a customized quadrotor simulator to conduct our experiments. For detailed formulations on quadrotor dynamics, we refer readers to [40]. To ensure a broad spectrum of domain randomization, we compute the inertia parameters based on the randomized size and mass of the quadrotor. This approach prevents the generation of unrealistic configurations by avoiding the independent randomization of these physical parameters.
We denote the arm length of the quadrotor as $L$. The width and height of the quadrotor are given by $\alpha_{w}L$ and $\alpha_{h}L$, respectively, where $\alpha_{w}$ and $\alpha_{h}$ represent the width-to-length and height-to-length ratios. Suppose the mass of the quadrotor is $M$, then the inertia matrix is computed as:
$\mathbf{I}=\frac{M}{12}\begin{pmatrix}L^{2}(\alpha_{w}^{2}+\alpha_{h}^{2})&0&0\\ 0&L^{2}(\alpha_{w}^{2}+\alpha_{h}^{2})&0\\ 0&0&L^{2}(2\alpha_{w}^{2})\end{pmatrix}$ (12)
B-A2 Dynamics Randomization
We randomize the parameters of the quadrotor dynamics for both training and evaluation, ensuring that the evaluation range is wider than the training range. For tasks with seen dynamics, we generate dynamics parameters within the training range. Conversely, for unseen dynamics, we use the evaluation range to generate the parameters to test the generalizablity of learned controllers.
TABLE V: Parameter ranges for dynamics randomization on quadrotor. $K_{p}^{0}=[37,37,11]$
. Parameter Name Train Range Eval Range Mass ($M$) [0.142, 0.950] [0.114, 1.140] Arm Length ($L$) [0.046, 0.200] [0.037, 0.240] Width-to-$L$ Ratio ($\alpha_{w}$) [1.0, 1.414] [0.852, 1.732] Height-to-$L$ Ratio ($\alpha_{h}$) [0.577, 1.0] [0.268, 1.414] Motor Force Constant ($K_{f}$) [1.15e-7, 7.64e-6] [9.16e-8, 9.17e-6] Max Thrust-to-Weight Ratio [2.0, 3.5] [1.6, 4.2] Motor Drag Constant ($\kappa$) [0.0041, 0.0168] [0.0033, 0.0201] Motor Time Constant ($\tau_{c}$) [0.3, 0.5] [0.24, 0.6] P Gains ($K_{p}$) [-0.3,0.3] ×$\times K_{p}^{0}$ [-0.5, 0.5] ×$\times K_{p}^{0}$
B-A3 Tasks
Here, we explain the tasks used in our experiments, which include one set-point task and four tracking tasks. Visualized samples for each trajectory are shown in Fig. 6, 7, 8 and 9. For all tasks, the controller frequency is set to 50 Hz, and each episode lasts for 10 seconds, i.e., 500 steps.
  • TrackCircle: This task requires the quadrotor to track a specific circular trajectory with a radius of 0.5 m and a period of 2 seconds. This trajectory is the simplest to track due to its smoothness and cyclicity.
  • TrackPoly: This task requires the quadrotor to track a random chained polynomial trajectory. Each trajectory consists of a series of random 5-degree polynomials. The starting point is $(x=0,y=0)$, and connection points between polynomials are randomly selected within the episode. Each polynomial lasts for a random duration between 1.5 and 4.0 seconds. We ensure the first, second, and third order derivatives are continuous at each point, resulting in a smooth trajectory.
  • TrackStar: This task requires the quadrotor to track a random five-pointed star trajectory. The center of the star is $(x=0,y=0)$, and the quadrotor starts from the center, flying to one point to track the star-shaped trajectory cyclically. The quadrotor is expected to fly at a certain velocity between points, making this task challenging due to its infeasibility.
  • TrackZigzag: This task requires the quadrotor to track a random zigzag trajectory. These trajectories are generated by randomly selecting time intervals between 1.0 and 2.0 seconds, generating waypoints after each interval, and linearly connecting each waypoint. The waypoints vary from -2 m to 2 m in the x, y, and z directions. This type of trajectory is the most challenging to track due to the highest randomness.
(a) Trajectory in X-Y plane.
(b) The first, second, and third order derivatives of the trajectory.
Figure 6: An example of the circular trajectory.
(a) Trajectory in X-Y plane.
(b) The first, second, and third order derivatives of the trajectory.
Figure 7: An example of the chained-polynomial trajectory.
(a) Trajectory in X-Y plane.
(b) The first and second order derivatives of the trajectory.
Figure 8: An example of the 5-point-star trajectory.
(a) Trajectory in X-Y plane.
(b) The first and second order derivatives of the trajectory.
Figure 9: An example of the zigzag trajectory.
B-B Quadrupedal Robots
For our experiments on quadrupedal robots, we implemented our method using the open-source code [^1] https://github.com/OpenRobotLab/HIMLoco of HIM [13]. We adhered to the hyper-parameters specified by HIM for conducting our experiments. Detailed information regarding the training parameters for reward functions, domain randomizations, and training terrains can be found in Appendix A of [13].
Training curriculum.
We adopt a training curriculum similar to that used in HIM, incorporating both terrain and command variations. Our height field map consists of 200 terrains arranged in a 20 × 10 grid, with each row featuring a specific type of terrain with increasing difficulty levels. Each grid cell measures 10 × 10 square meters. To facilitate the training process, robots initially start on terrains with the lowest difficulty. As the robots successfully traverse more than half of the grid, we incrementally increase the terrain difficulty. Conversely, robots that fail to cover at least half of the required distance are moved to simpler terrains.
[…]
Evaluation tasks.
[…]
Appendix C PPO Hyper-parameters
For experiments on quadrotors, we use the identical hyper-parameters in PPO training for all controllers.
TABLE VI: Hyper Parameters for Training Controllers for Quadrotors.
Hyper Parameters for Training
Batch Size $1024\times 128$
Mini-batch Size 1024
Number of Epochs 5
Clip Range 0.2
Entropy Coefficient 0.01
Discount Factor 0.99
GAE Discount Factor 0.95
Desired KL-Divergence 0.01
Learning Rate $1\times 10^{-3}$
Adam Epsilon $1\times 10^{-8}$
Gradient Clipping 10
For experiments on quadrupedal, we follow the hyper-parameter settings from [13], while our methods including one variant based on HIM share the same hyper-parameters.
TABLE VII: Hyper Parameters for Training Controllers for Quadrupedal Robots.
Hyper Parameters for Training
General
Batch Size $4096\times 200$
Mini-batch Size $4096\times 50$
Number of Epochs 5
PPO
Clip Range 0.2
Entropy Coefficient 0.01
Discount Factor 0.99
GAE Discount Factor 0.95
Desired KL-Divergence 0.01
Learning Rate $1\times 10^{-3}$
Adam Epsilon $1\times 10^{-8}$
Gradient Clipping 10
HIM
Number of Prototypes 16
Contrastive Loss Scale 1.0
Velocity Estimation Loss Scale 1.0
Learning Rate $1\times 10^{-3}$
Adam Epsilon $1\times 10^{-8}$
Gradient Clipping 10

### v2

@@
 Figure 2: The closed-loop NeuralIMC framework using an explicit predictive model to calculate the next body state, applirope to both quadrotors and legged robots.
-In this paper, we propose a general and flexible framework that integrates model-based and RL-based control.
+In this paper, we propose a new learning-based control methods which leverages a predictive error feedback to enhance control robustness, inspired by model-based control.
 By simplifying the model to focus on the robot’s main body and treating its dynamics as a rigid transformation, we bypass the high-dimensional nonlinearity of complex robot systems and environmental interactions.
@@
 However, applying IMC to more complex and nonlinear systems is challenging due to the difficulty of crafting reliable models for such systems.
-Long et al.
-[13] further proposed Hybrid Internal Model, in which they leverage a learning-based internal model with unsupervised contrastive learning.
-The learned model serves as a predictive encoder for estimating the system responses in the latent space.
+Hybrid Internal Model (HIM) further leverages a learning-based internal model that employs unsupervised contrastive learning, with the learned model acting as a predictive encoder for estimating system responses in the latent space, as proposed by Long et al.
+[13].
 Despite their advantages, these methods either rely on explicit disturbance estimation or are heavily influenced by prediction accuracy, which can significantly affect performance if compromised.
@@
 Conversely, DATT [20] incorporates an $\mathcal{L}_{1}$ adaptation law [21] to estimate forces applied to quadrotors directly, which are then used as policy inputs.
-Despite its effectiveness with precise parameter tuning, this method is sensitive to handcrafted settings and challenging to generalize to complex systems.
+CAL [28] integrates RL with nonlinear model predictive control (NMPC) by employing RL to generate residual actions that augment NMPC’s outputs.
+Although both DATT and CAL have proven effective with precise parameter tuning, they remain sensitive to handcrafted settings in their classical control components, limiting their generalizability to more complex systems.
 III Methods
@@
 Figure 3: Feedback structures of classical feedback control and Internal Model Control (IMC).
-Internal Model Control (IMC) is a powerful structure for robust control, offering significant advantages over classical feedback control, particularly in managing model uncertainties and external disturbances.
-Unlike classical feedback control, which relies on real-time error correction based on the difference between desired and actual outputs, IMC incorporates a predictive model of the process, allowing it to anticipate and compensate for disturbances more effectively.
-The structures of classical feedback control and IMC are illustrated in Fig.
-3, where $P$ and $\tilde{P}$ represent the plant and internal models, respectively.
-$Q$ is the IMC controller, and $C$ is the classical feedback controller.
-The reference signal, disturbance, and system output are denoted by $r$, $d$, and $y$, respectively.
-In IMC, the predicted response $\tilde{y}$ is subtracted from the actual response $y$, and the difference, $\tilde{d}=\tilde{y}-y$, is used as the input signal to the IMC controller.
-Generally, $\tilde{d}$ is non-zero due to modeling errors and unknown disturbances not accounted for in the model.
-Consequently, IMC can provide more robust performance by continuously adjusting the control input according to the discrepancy between the predicted and actual responses.
-While the design procedure for IMC is straightforward when the plant system is easy to model [3], it becomes significantly more complex and challenging for systems with nonlinear dynamics or when accurate models are difficult to obtain [30, 31].
-This complexity arises because methods like linearization or approximation often fail to capture the true behavior of nonlinear systems, leading to performance degradation.
-Such a dependency on precise modeling limits its practical applications in robotic systems, such as quadrotors with cable-suspended payloads and legged robots on complex terrains, where achieving a robust and accurate system model is inherently challenging.
-To address these challenges, our method combines the strengths of the IMC structure with large-scale RL training.
-A simplified predictive model in IMC provides a rough but stable estimate of the system’s response.
-Instead of relying solely on precise system modeling, our approach uses large-scale randomized simulations to learn a robust control policy that compensates for inaccuracies and simplified-out complexities.
-Through extensive training, the RL algorithm optimizes the control policy conditioned on predictive errors, enhancing robustness and performance in real-world scenarios where exact modeling is impractical.
-The predictive error reveals signals about external disturbances and internal characteristics of nonlinear dynamics, enabling more informative control decisions and effective policy learning.
+Internal Model Control (IMC) is a robust control framework that outperforms classical feedback control in handling model uncertainties and disturbances.
+Unlike feedback control, which corrects errors in real time, IMC uses a predictive model to anticipate and compensate for disturbances.
+Fig.
+3 illustrates their structures, where $P$ and $\tilde{P}$ are the plant and internal models, $Q$ is the IMC controller, and $C$ is the classical feedback controller.
+The reference, disturbance, and output are denoted by $r$, $d$, and $y$, respectively.
+IMC computes the difference $\tilde{d}=\tilde{y}-y$ between predicted and actual responses, adjusting control inputs accordingly for enhanced robustness.
+IMC is straightforward when the plant is easily modeled [3] but becomes challenging for nonlinear systems where approximations fail to capture true dynamics [30, 31].
+This reliance on precise models limits its use in robotics, such as quadrotors with suspended payloads or legged robots on uneven terrain, where accurate system modeling is difficult.
+Our method integrates IMC with large-scale RL training to address these challenges.
+A simplified predictive model provides a rough yet stable estimate, while large-scale randomized simulations train a robust control policy that compensates for modeling inaccuracies.
+RL optimizes the policy with the predictive error feedback, revealing external disturbances and nonlinear dynamics, leading to more adaptive and effective control in real-world settings.
 III-B Design of a Simplified Predictive Model
-The predictive model is crucial in our framework, providing the baseline needed to detect and compensate for system discrepancies and disturbances.
-Designing an effective predictive model involves two main challenges: accurately estimating the system response and ensuring the model’s predictions closely match actual system behavior.
-Accurate estimation is essential for the feedback loop to reflect the true robot state, while alignment between predictions and real responses is critical for reliable performance under varying conditions.
+The predictive model is key to our framework, enabling discrepancy detection and compensation.
+Its design faces two challenges: accurately estimating system response and closely aligning predictions with actual behavior.
+Precise estimation ensures the feedback loop reflects the true robot state, while alignment ensures reliability under varying conditions.
 To streamline state estimation across diverse robots, we model only the rigid body state, excluding actuator- or limb-specific states.
@@
 Finally, we can predict the next quaternion and angular velocity as follows, where $\otimes$ denotes quaternion multiplication:
-$\tilde{\mathbf{q}}_{k+1}$ $=\mathbf{q}_{k}\otimes\mathbf{q}_{\Delta\theta},$ (5)
-$\tilde{\boldsymbol{\omega}}_{k+1}$ $=\boldsymbol{\omega}_{k}+\dot{\boldsymbol{\omega}}_{k}\Delta t.$ (6)
+$\tilde{\mathbf{q}}_{k+1}$ $=\mathbf{q}_{k}\otimes\mathbf{q}_{\Delta\theta}$ (5)
+$\tilde{\mathbf{\omega}}_{k+1}$ $=\mathbf{\omega}_{k}+\dot{\mathbf{\omega}}_{k}\Delta t$ (6)
 We have established an explicit method to predict the next body state, assuming the required motion can be represented by commands at the acceleration or velocity level.
-Then, given the predicted body state $\tilde{\mathbf{x}}_{k+1}=\left[\tilde{\mathbf{p}}_{k+1},\tilde{\mathbf{q}}_{k+1},\tilde{\mathbf{v}}_{k+1},\tilde{\boldsymbol{\omega}}_{k+1}\right]$ and the actual body state $\mathbf{x}_{k+1}$, we compute the predictive error $\tilde{\mathbf{d}}_{k+1}$, which is then fed into the policy to provide effective feedback signals.
-$\tilde{\mathbf{d}}_{k+1}=\left[\Delta\tilde{\mathbf{p}}_{k+1},1-\cos{\tilde{\Theta}},\Delta\tilde{\mathbf{v}}_{k+1},\Delta\tilde{\boldsymbol{\omega}}_{k+1}\right]$ (7)
-Here, $\Delta\tilde{\mathbf{p}}_{k+1}$, $\Delta\tilde{\mathbf{v}}_{k+1}$, and $\Delta\tilde{\boldsymbol{\omega}}_{k+1}$ represent the differences in position, velocity, and angular velocity, respectively.
+Then, given the predicted body state $\tilde{\mathbf{x}}_{k+1}=\left[\tilde{\mathbf{p}}_{k+1},\tilde{\mathbf{q}}_{k+1},\tilde{\mathbf{v}}_{k+1},\tilde{\mathbf{\omega}}_{k+1}\right]$ and the actual body state $\mathbf{x}_{k+1}$, we compute the predictive error $\tilde{\mathbf{d}}_{k+1}$, which is then fed into the policy to provide effective feedback signals.
+$\tilde{\mathbf{d}}_{k+1}=\left[\Delta\tilde{\mathbf{p}}_{k+1},1-\cos{\tilde{\Theta}},\Delta\tilde{\mathbf{v}}_{k+1},\Delta\tilde{\mathbf{\omega}}_{k+1}\right]$ (7)
+Here, $\Delta\tilde{\mathbf{p}}_{k+1}$, $\Delta\tilde{\mathbf{v}}_{k+1}$, and $\Delta\tilde{\mathbf{\omega}}_{k+1}$ represent the differences in position, velocity, and angular velocity, respectively.
 We use the cosine distance to quantify the difference between orientations, where $\tilde{\Theta}$ is the angle difference between the two quaternions.
@@
 III-C1 Quadrotors
-Traditional quadrotor controllers often use multi-layered cascaded structures [38], offering flexible options for integrating neural controllers.
-Recent research shows that using collective thrust and bodyrate (CTBR) as the action space for neural controllers provides significant advantages [39, 20, 9].
-CTBR-based quadrotor controllers exhibit strong resilience to dynamics mismatches, effective domain transfer, and high agility and performance.
-This action space design is inherently compatible with our framework.
-We can directly use the bodyrate command from the policy to update the rotation, while the linear acceleration is computed using the collective thrust to translate the robot.
-Consequently, the predictive error, calculated from the previous policy output and the latest successive body states, can be seamlessly integrated into the current policy input without further modifications to the learning algorithm.
+Traditional quadrotor controllers use cascaded structures [38], allowing flexible integration of neural controllers.
+Recent studies show that using collective thrust and bodyrate (CTBR) as the action space enhances resilience to dynamics mismatches, improves domain transfer, and boosts agility [39, 20, 9].
+This design aligns with our framework: the policy’s bodyrate command updates rotation, while collective thrust determines linear acceleration.
+Consequently, predictive errors, computed from prior policy output and successive body states, integrate directly into the current policy input without modifying the learning algorithm.
 III-C2 Quadrupedal robots
-Unlike quadrotors, legged robots, particularly quadrupedal and bipedal robots, have garnered significant attention for RL-based controllers due to their higher control dimensions and the complexities of modeling interactions with diverse terrains [8, 11, 12, 14, 19].
-Typically, a neural controller for legged robots generates target actuator positions to achieve the desired body velocity, including forward velocity ($v_{x}$), lateral velocity ($v_{y}$), and yaw rate ($\omega_{z}$).
-To integrate our predictive model into the control loop, we use the reference velocity command as the model input instead of the high-dimensional actuator motions.
-By differentiating two consecutive reference velocity commands, we obtain the linear acceleration in x and y axes, and the angular acceleration for the yaw angle.
-In this way, we can apply the rigid transformation by setting other acceleration terms to zero, ensuring seamless integration into the existing policy structure.
-This method is compatible with other algorithms by simply adding a plug-and-play feedback loop to augment policy inputs.
+Unlike quadrotors, legged robots, particularly quadrupeds and bipeds, are well-suited for RL-based control due to their high-dimensional actuation and terrain interactions [8, 11, 12, 14, 19].
+Neural controllers typically generate actuator targets to achieve desired body velocities ($v_{x},v_{y},\omega_{z}$).
+To integrate our predictive model into the control loop,, we use reference velocity commands as model inputs, deriving linear and angular accelerations by differentiating consecutive commands.
+Setting other accelerations to zero ensures seamless integration with existing policies.
+This method is compatible with other algorithms via a plug-and-play feedback loop augmenting policy inputs.
 IV Experimental Results
@@
 Bold values denote the best results.
-Blue cells indicate performance under random wind fields, with percentages in brackets representing the relative change in error compared to static conditions.
+Blue cells indicate performance under random external force disturbances, with percentages in brackets representing the relative change in error compared to static conditions.
 T and S refer to the teacher and student models for RMA, respectively, while DATT with GT indicates the use of ground-truth disturbance inputs instead of $\mathcal{L}_{1}$ estimates.
@@
 Circle Chained Poly 5-Point Star Zigzag
-Train MPPI+RL 0.087 0.157 (+80.46%) 0.056 0.118 (+110.71%) 0.103 0.175 (+69.90%) 0.105 0.164 (+56.19%)
+Train $\mathcal{L}_{1}$-PID 0.214 0.237 (+10.74%) 0.043 0.054 (+25.58%) crashed crashed
+$\mathcal{L}_{1}$-MPPI 0.202 0.215 (+6.44%) 0.092 0.117 (+27.17%) 0.193 0.216 (+11.91%) 0.169 0.186 (+10.06%)
+CAL 0.087 0.157 (+80.46%) 0.056 0.118 (+110.71%) 0.103 0.175 (+69.90%) 0.105 0.164 (+56.19%)
 PPO 0.102 0.169 (+65.69%) 0.065 0.117 (+80.00%) 0.097 0.172 (+77.32%) 0.095 0.154 (+62.11%)
@@
 Ours 0.072 0.083 (+15.28%) 0.032 0.038 (+18.75%) 0.065 0.073 (+12.31%) 0.057 0.064 (+12.28%)
-Eval MPPI+RL 0.224 0.321 (+43.30%) 0.171 0.251 (+46.78%) 0.262 0.364 (+38.93%) 0.265 0.358 (+35.09%)
+Eval $\mathcal{L}_{1}$-PID 0.372 0.416 (+11.82%) 0.061 0.107 (+75.41%) crashed crashed
+$\mathcal{L}_{1}$-MPPI 0.212 0.234 (+10.38%) 0.097 0.132 (+36.08%) 0.207 0.253 (+22.22%) 0.176 0.218 (+23.86%)
+CAL 0.224 0.321 (+43.30%) 0.171 0.251 (+46.78%) 0.262 0.364 (+38.93%) 0.265 0.358 (+35.09%)
 PPO 0.265 0.351 (+32.45%) 0.183 0.273 (+49.18%) 0.257 0.374 (+45.53%) 0.265 0.367 (+38.49%)
@@
 IV-B Experiments on Quadrotors
-We compare our method for quadrotors against four baselines: MPPI+RL [28], PPO [27], RMA [12], DATT [20], using two variants each for RMA and DATT to enable deeper analyses.
+We compare our method for quadrotors against two classical control methods ($\mathcal{L}_{1}$-PID and $\mathcal{L}_{1}$-MPPI) and four learning-based baselines (CAL [28], PPO [27], RMA [12], DATT [20]), using two variants each for RMA and DATT to enable deeper analyses.
 All methods use the same PPO training parameters.
-For MPPI+RL, MPPI serves as a sampling-based nonlinear MPC method, with the RL policy compensating the MPPI outputs.
+For CAL [28], a sum combination of MPPI and RL, MPPI serves as a sampling-based non-linear MPC method, with the RL policy compensating the MPPI output.
 We set the sampling number to 4096 and the rollout horizon to 20 in MPPI’s sampling procedure.
@@
 II, our method consistently outperforms others across nearly all tasks, particularly under disturbances and out-of-distribution system dynamics.
+Compared to state-of-the-art learning-based methods, classical adaptive control methods perform worse due to their reliance on handcrafted parameters, which fail to generalize across diverse conditions.
 Among the baselines, DATT shows strong resilience to external disturbances, supported by its $\mathcal{L}_{1}$ estimator.
@@
 None PPO 0.132 0.085 0.143 0.087 0.200 0.085 0.163 0.085
+RMA 0.084 0.050 0.086 0.051 0.127 0.050 0.110 0.050
 HIM 0.100 0.067 0.103 0.068 0.181 0.067 0.147 0.067
@@
 Impulse PPO 0.151 0.101 0.162 0.103 0.216 0.101 0.178 0.101
+RMA 0.096 0.059 0.098 0.060 0.137 0.060 0.120 0.059
 HIM 0.117 0.080 0.120 0.081 0.186 0.080 0.157 0.080
@@
 IV-C Experiments on Quadrupedal Robots
-For quadrupedal robots, we compare our methods with PPO [27] and HIM [13].
-PPO serves as a standard baseline in continuous control for quadrupeds, while HIM leverages unsupervised contrastive learning to train a predictive encoder for the next latent state and body velocity.
+For quadrupedal robots, we compare our methods with PPO [27], RMA [12] and HIM [13].
+PPO and RMA serve as standard and common baselines in continuous control for quadrupeds, while HIM leverages unsupervised contrastive learning to train a predictive encoder for the next latent state and body velocity.
+Here, we directly use the teacher-policy of RMA to set a strong baseline.
 Our method enhances the PPO policy by integrating predictive error feedback.
@@
 As shown in Tab.
-III, our method demonstrates superior performance, achieving relative improvements over HIM ranging from approximately 26% to 44% across various test conditions.
+III, our method outperforms HIM by 26–44% across various test conditions and matches or surpasses the privileged RMA.
 The combination of our framework with HIM also shows improvements over the original HIM.
@@
 IV-D Ablation Studies
-In this section, we conduct two ablation studies to further explore the proposed framework.
-First, inspired by [14], we investigate whether integrating short- and long-horizon input/output histories into the feedback structure can enhance performance.
-Next, we introduce noise into the internal predictive model to test its noise tolerance and assess the impact of prediction accuracy on control performance.
-We perform all experiments on quadrotors tracking zigzag trajectories across three random seeds.
+In this section, we conduct three ablation studies to analyze our framework.
+Inspired by [14], we first examine whether incorporating short- and long-horizon input/output histories improves performance.
+Next, we introduce noise into the predictive model’s input and output to assess its robustness and the impact of prediction accuracy on control.
+Finally, we compare our method using the predictive model versus a full-dynamics model to evaluate the effect of model simplification.
+All experiments involve quadrotors tracking zigzag trajectories across three random seeds.
 (a) Varying short-history length.
@@
 Noise $\sigma_{n}$ Train Eval
-static windy static windy
+static disturbed static disturbed
 0. 0.057 0.064 (+12.28%) 0.101 0.119 (+17.82%)
-0.02 0.068 0.077 (+13.24%) 0.136 0.172 (+26.47%)
-0.04 0.077 0.103 (+33.77%) 0.161 0.230 (+42.86%)
-0.08 0.083 0.130 (+56.63%) 0.201 0.285 (+41.79%)
-0.16 0.087 0.150 (+72.41%) 0.240 0.343 (+42.92%)
+input noise 0.02 0.063 0.069 (+9.52%) 0.092 0.117 (+27.17%)
+0.04 0.070 0.094 (+34.28%) 0.100 0.154 (+54.00%)
+0.08 0.071 0.121 (+70.42%) 0.134 0.202 (+50.74%)
+0.16 0.082 0.135 (+64.63%) 0.152 0.231 (+51.97%)
+output noise 0.02 0.063 0.071 (+12.69%) 0.076 0.132 (+73.68%)
+0.04 0.073 0.097 (+32.87%) 0.102 0.161 (+57.84%)
+0.08 0.074 0.121 (+63.51%) 0.118 0.179 (+51.69%)
+0.16 0.076 0.130 (+71.05%) 0.138 0.222 (+60.86%)
 (a) Quadrotor system for deployment.
@@
 IV-D1 History structure
-We evaluate different feedback structures incorporating historical records.
-In short-history experiments, we concatenate several recent input/output pairs, including predictive error within each pair, into a single input vector for the policy.
-For long-history experiments, we retain the latest state-action pair while providing additional historical context.
-Following prior work [12, 14], we utilize a TCN model to process long-horizon history into a latent vector, which is then concatenated with the current input for the policy.
+We evaluate feedback structures incorporating historical records.
+In short-history experiments, we concatenate recent input/output pairs, including predictive errors, into a single input vector.
+For long-history experiments, we retain the latest state-action pair while encoding additional history using a TCN model [12, 14], concatenating its latent output with the policy input.
 As shown in Fig.
-4, increasing the length of historical inputs in the short- or long-horizon period does not improve performance compared to using the immediate response (i.e., using only the lated one-step short history).
-Although this differs from [14], it aligns with classical adaptive control methods such as $\mathcal{L}_{1}$ adaptive control [21] and robust control methods like internal model control [2, 3], where immediate error feedback suffices for improving motion accuracy and robustness.
-This simpler structure eases implementation and avoids cumulative error issues often encountered in real robot systems.
+4, using longer history—either short- or long-term—does not improve performance over immediate feedback (i.e., the latest one-step history).
+While differing from [14], this aligns with $\mathcal{L}_{1}$ adaptive control [21] and robust control methods [2, 3], where immediate error feedback suffices for accuracy and robustness.
+This simpler structure also reduces implementation complexity and mitigates cumulative errors in real robots.
 IV-D2 Noise tolerance
-To assess the impact of the internal predictive model’s accuracy on control performance, we introduce varying levels of noise into the model’s predictions.
-Specifically, we add zero-mean Gaussian noise with different variances to the responses of our simplified internal model.
+To assess the impact of the internal predictive model’s accuracy on control performance, we introduce varying levels of noise into the model’s state input or predictions.
+Specifically, we introduce zero-mean Gaussian noise with varying variances to the state vector, including position, velocity, quaternion, and angular velocity.
+For the quaternion, we treat the sampled Gaussian noise as a delta Euler angle, convert it to a delta quaternion, and apply it to rotate the original orientation.
 We train and evaluate multiple policies using a range of noisy predictive models for quadrotor tracking on zigzag trajectories, with the results presented in Tab.
 IV.
-The results indicate that our method maintains competitive performance under slight noise, though the accuracy of the predictive model becomes increasingly critical for larger noise levels, particularly in scenarios with external disturbances or out-of-distribution system dynamics.
-Across different training and evaluation ranges, performance declines as noise intensifies, with a similar trend observed under windy force disturbances.
+The results show that our method maintains competitive performance under mild noise, whether applied to model inputs or responses.
+However, as noise levels increase, the accuracy of the predictive model becomes increasingly crucial, particularly in scenarios with external disturbances or out-of-distribution system dynamics.
+Across different training and evaluation ranges, performance declines as noise intensifies, with a similar trend observed under external force disturbances.
 This observation highlights an advantage of our explicit predictive model over other learning-based approaches: it provides a stable baseline estimation, albeit approximate.
@@
 While this model simplifies nonlinear complexities, the RL part in our framework compensates effectively within an informative closed-loop system.
+IV-D3 Model simplification
+TABLE V: Evaluation of the internal predictive model’s complexity, tested on quadrotors tracking zigzag trajectories.
+Model Train Eval
+static disturbed static disturbed
+Simplified 0.057 0.064 (+12.28%) 0.101 0.119 (+17.82%)
+Full 0.061 0.067 (+9.84%) 0.095 0.105 (+10.53%)
+We investigate the impact of predictive model simplification on performance.
+A quadrotor serves as our testbed due to its well-studied dynamics model [40].
+We compare our approach with a variant that uses the full-dynamics model for tracking zigzag trajectories.
+As shown in Tab.
+V, the results indicate that despite the simplification, performance degradation is marginal.
+Notably, this simplification improves generalization across different morphologies, making it applicable to complex systems that are difficult to model.
 IV-E Real-world Deployment
@@
 V Conclusion
-In this paper, we propose a novel framework, Neural Internal Model Control (NeuralIMC), that combines model-based control with RL-based neural control.
-By focusing on body dynamics and adopting the white-box rigid-body transformation, we bypass high-dimensional nonlinearity and employ a simplified explicit predictive model to form a closed-loop feedback structure.
-Integrating model-free RL algorithms allows the neural control policy to react according to the predictive error feedback.
-Our method demonstrates superior performance on quadrotors and quadrupedal robots, surpassing state-of-the-art control methods in both domains.
-Real-world deployment on a quadrotor with heavy payloads highlights the robustness of our approach in handling disturbances and adapting to varying dynamics.
-This work underscores the potential of our combined framework for enhanced robustness and adaptability in robotic systems and suggests a direction for more interpretable learning-based control systems in future research.
-Limitations.
-While extensive simulation experiments highlight the advantages of the proposed neural control framework, real-world evaluation remains limited.
-We plan to test our framework on a wider range of robots, including quadrotors of varying sizes and legged robots on complex terrains.
-Additionally, our method can be further strengthened by incorporating more effective disturbance training [45].
-Appendix A Specification Details
-In this section, we describe the integration of the simplified predictive model discussed in Sec.
-III-B into the neural control pipeline of two widely used robotic systems: quadrotors and quadrupedal robots.
-We specifically detail the process of obtaining the necessary model inputs to update the predicted body states.
-For clarity, we omit the subscript $k$; all states—position $\mathbf{x}$, quaternion $\mathbf{q}$, velocity $\mathbf{v}$, and angular velocity $\boldsymbol{\omega}$—are from the previous time step, i.e., $k=t-1$.
-Our goal is to predict the expected body state at the current time step $t$.
-A-A Quadrotors
-For quadrotors, as demonstrated in Sec.
-III-C, we define the action space of the policy using mass-normalized collective thrust and body rates (CTBR).
-This is followed by a PD controller to manage the quadrotor’s flight dynamics.
-We denote the collective thrust as $\mathbf{a}_{z}^{\mathcal{B}}$ and the body rates as $\boldsymbol{\omega}^{\mathcal{B}}$, both expressed in the quadrotor’s body frame.
-The acceleration in the world frame can then be computed as follows:
-$\mathbf{a}=\mathcal{T}_{\mathbf{q}}(\mathbf{a}_{z}^{\mathcal{B}})+\mathbf{g}$ (8)
-where $\mathcal{T}_{\mathbf{q}}$ represents the rotation of the z-axis acceleration from the body frame to the world frame, and $\mathbf{g}=[0,0,-9.81]^{T}$ is the free-fall acceleration.
-Then, we can apply Eqs.
-1 and 2 to obtain the predicted position $\tilde{\mathbf{x}}$ and velocity $\tilde{\mathbf{v}}$.
-For the rotation, we set $\dot{\boldsymbol{\omega}}$ to zero and directly replace $\boldsymbol{\omega}$ with $\boldsymbol{\omega}^{\mathcal{B}}$.
-Then, we can apply Eqs.
-3, 4, 5 and 6 to obtain the predicted quaternion $\tilde{\mathbf{q}}$ and angular velocity $\tilde{\boldsymbol{\omega}}$.
-A-B Quadrupedal Robots
-For quadrupedal robots, we utilize reference velocity commands as the model inputs rather than the policy outputs.
-Given that the provided velocity commands do not cover all degrees of freedom, we set the unspecified parts to zero.
-We then use the command values as target quantities, calculating the required acceleration by taking the difference between the target and the actual state of the robot, in order to update the state.
-Given the reference forward velocity $\hat{v}_{x}^{\mathcal{B}}$, lateral velocity $\hat{v}_{y}^{\mathcal{B}}$, and yaw rate $\hat{\omega}_{z}^{\mathcal{B}}$, all defined in the body frame, we can compute the desired acceleration as follows:
-$\mathbf{a}^{\mathcal{B}}=\begin{bmatrix}(\hat{v}_{x}^{\mathcal{B}}-v_{x}^{\mathcal{B}})/{\Delta t}\\ (\hat{v}_{y}^{\mathcal{B}}-v_{y}^{\mathcal{B}})/{\Delta t}\\ 0\end{bmatrix}$ (9)
-In this context, $v_{x}^{\mathcal{B}}$ and $v_{y}^{\mathcal{B}}$ denote the forward and lateral velocities of the body from the previous step, respectively, transformed into the body frame.
-The world-frame acceleration can then be obtained by applying the transformation $\mathcal{T}_{\mathbf{q}}$:
-$\mathbf{a}=\mathcal{T}_{\mathbf{q}}(\mathbf{a}^{\mathcal{B}})$ (10)
-Similarly, the angular acceleration can be computed as:
-$\dot{\boldsymbol{\omega}}=\begin{bmatrix}0\\ 0\\ \hat{\omega}_{z}^{\mathcal{B}}-\omega_{z}^{\mathcal{B}}\end{bmatrix}$ (11)
-Subsequently, the computed acceleration $\mathbf{a}$ and angular acceleration $\dot{\boldsymbol{\omega}}$ can be substituted into the relevant equations (Eqs.
-1, 2, 3, 4, 5 and 6) to predict the next body state $\tilde{\mathbf{x}}$.
-Appendix B Experiment Details
-In this section, we provide a detailed explanation of how we set up the simulation tasks for training and evaluating various controllers for both quadrotors and quadrupedal robots.
-B-A Quadrotors
-B-A1 Quadrotor Dynamics
-We employ a customized quadrotor simulator to conduct our experiments.
-For detailed formulations on quadrotor dynamics, we refer readers to [40].
-To ensure a broad spectrum of domain randomization, we compute the inertia parameters based on the randomized size and mass of the quadrotor.
-This approach prevents the generation of unrealistic configurations by avoiding the independent randomization of these physical parameters.
-We denote the arm length of the quadrotor as $L$.
-The width and height of the quadrotor are given by $\alpha_{w}L$ and $\alpha_{h}L$, respectively, where $\alpha_{w}$ and $\alpha_{h}$ represent the width-to-length and height-to-length ratios.
-Suppose the mass of the quadrotor is $M$, then the inertia matrix is computed as:
-$\mathbf{I}=\frac{M}{12}\begin{pmatrix}L^{2}(\alpha_{w}^{2}+\alpha_{h}^{2})&0&0\\ 0&L^{2}(\alpha_{w}^{2}+\alpha_{h}^{2})&0\\ 0&0&L^{2}(2\alpha_{w}^{2})\end{pmatrix}$ (12)
-B-A2 Dynamics Randomization
-We randomize the parameters of the quadrotor dynamics for both training and evaluation, ensuring that the evaluation range is wider than the training range.
-For tasks with seen dynamics, we generate dynamics parameters within the training range.
-Conversely, for unseen dynamics, we use the evaluation range to generate the parameters to test the generalizablity of learned controllers.
-TABLE V: Parameter ranges for dynamics randomization on quadrotor. $K_{p}^{0}=[37,37,11]$
-.
-Parameter Name Train Range Eval Range Mass ($M$) [0.142, 0.950] [0.114, 1.140] Arm Length ($L$) [0.046, 0.200] [0.037, 0.240] Width-to-$L$ Ratio ($\alpha_{w}$) [1.0, 1.414] [0.852, 1.732] Height-to-$L$ Ratio ($\alpha_{h}$) [0.577, 1.0] [0.268, 1.414] Motor Force Constant ($K_{f}$) [1.15e-7, 7.64e-6] [9.16e-8, 9.17e-6] Max Thrust-to-Weight Ratio [2.0, 3.5] [1.6, 4.2] Motor Drag Constant ($\kappa$) [0.0041, 0.0168] [0.0033, 0.0201] Motor Time Constant ($\tau_{c}$) [0.3, 0.5] [0.24, 0.6] P Gains ($K_{p}$) [-0.3,0.3] ×$\times K_{p}^{0}$ [-0.5, 0.5] ×$\times K_{p}^{0}$
-B-A3 Tasks
-Here, we explain the tasks used in our experiments, which include one set-point task and four tracking tasks.
-Visualized samples for each trajectory are shown in Fig.
-6, 7, 8 and 9.
-For all tasks, the controller frequency is set to 50 Hz, and each episode lasts for 10 seconds, i.e., 500 steps.
-  • TrackCircle: This task requires the quadrotor to track a specific circular trajectory with a radius of 0.5 m and a period of 2 seconds.
-This trajectory is the simplest to track due to its smoothness and cyclicity.
-  • TrackPoly: This task requires the quadrotor to track a random chained polynomial trajectory.
-Each trajectory consists of a series of random 5-degree polynomials.
-The starting point is $(x=0,y=0)$, and connection points between polynomials are randomly selected within the episode.
-Each polynomial lasts for a random duration between 1.5 and 4.0 seconds.
-We ensure the first, second, and third order derivatives are continuous at each point, resulting in a smooth trajectory.
-  • TrackStar: This task requires the quadrotor to track a random five-pointed star trajectory.
-The center of the star is $(x=0,y=0)$, and the quadrotor starts from the center, flying to one point to track the star-shaped trajectory cyclically.
-The quadrotor is expected to fly at a certain velocity between points, making this task challenging due to its infeasibility.
-  • TrackZigzag: This task requires the quadrotor to track a random zigzag trajectory.
-These trajectories are generated by randomly selecting time intervals between 1.0 and 2.0 seconds, generating waypoints after each interval, and linearly connecting each waypoint.
-The waypoints vary from -2 m to 2 m in the x, y, and z directions.
-This type of trajectory is the most challenging to track due to the highest randomness.
-(a) Trajectory in X-Y plane.
-(b) The first, second, and third order derivatives of the trajectory.
-Figure 6: An example of the circular trajectory.
-(a) Trajectory in X-Y plane.
-(b) The first, second, and third order derivatives of the trajectory.
-Figure 7: An example of the chained-polynomial trajectory.
-(a) Trajectory in X-Y plane.
-(b) The first and second order derivatives of the trajectory.
-Figure 8: An example of the 5-point-star trajectory.
-(a) Trajectory in X-Y plane.
-(b) The first and second order derivatives of the trajectory.
-Figure 9: An example of the zigzag trajectory.
-B-B Quadrupedal Robots
-For our experiments on quadrupedal robots, we implemented our method using the open-source code [^1] https://github.com/OpenRobotLab/HIMLoco of HIM [13].
-We adhered to the hyper-parameters specified by HIM for conducting our experiments.
-Detailed information regarding the training parameters for reward functions, domain randomizations, and training terrains can be found in Appendix A of [13].
-Training curriculum.
-We adopt a training curriculum similar to that used in HIM, incorporating both terrain and command variations.
-Our height field map consists of 200 terrains arranged in a 20 × 10 grid, with each row featuring a specific type of terrain with increasing difficulty levels.
-Each grid cell measures 10 × 10 square meters.
-To facilitate the training process, robots initially start on terrains with the lowest difficulty.
-As the robots successfully traverse more than half of the grid, we incrementally increase the terrain difficulty.
-Conversely, robots that fail to cover at least half of the required distance are moved to simpler terrains.
-[…]
-Evaluation tasks.
-[…]
-Appendix C PPO Hyper-parameters
-For experiments on quadrotors, we use the identical hyper-parameters in PPO training for all controllers.
-TABLE VI: Hyper Parameters for Training Controllers for Quadrotors.
-Hyper Parameters for Training
-Batch Size $1024\times 128$
-Mini-batch Size 1024
-Number of Epochs 5
-Clip Range 0.2
-Entropy Coefficient 0.01
-Discount Factor 0.99
-GAE Discount Factor 0.95
-Desired KL-Divergence 0.01
-Learning Rate $1\times 10^{-3}$
-Adam Epsilon $1\times 10^{-8}$
-Gradient Clipping 10
-For experiments on quadrupedal, we follow the hyper-parameter settings from [13], while our methods including one variant based on HIM share the same hyper-parameters.
-TABLE VII: Hyper Parameters for Training Controllers for Quadrupedal Robots.
-Hyper Parameters for Training
-General
-Batch Size $4096\times 200$
-Mini-batch Size $4096\times 50$
-Number of Epochs 5
-PPO
-Clip Range 0.2
-Entropy Coefficient 0.01
-Discount Factor 0.99
-GAE Discount Factor 0.95
-Desired KL-Divergence 0.01
-Learning Rate $1\times 10^{-3}$
-Adam Epsilon $1\times 10^{-8}$
-Gradient Clipping 10
-HIM
-Number of Prototypes 16
-Contrastive Loss Scale 1.0
-Velocity Estimation Loss Scale 1.0
-Learning Rate $1\times 10^{-3}$
-Adam Epsilon $1\times 10^{-8}$
-Gradient Clipping 10
+In this paper, we propose Neural Internal Model Control (NeuralIMC), a framework that bridges model-based control with RL-based neural control.
+By leveraging body dynamics and a white-box rigid-body transformation, we avoid high-dimensional nonlinearity, using a simplified predictive model in a closed-loop feedback structure.
+Model-free RL enables the policy to react to predictive errors.
+NeuralIMC outperforms state-of-the-art controllers on quadrotors and quadrupedal robots.
+Real-world tests on a heavily loaded quadrotor demonstrate robustness to disturbances and dynamic variations, highlighting its potential for interpretable and adaptive learning-based control.
+For the future work, we’d like to conduct more expanded real-world experiments includes testing on diverse robots and improving disturbance training [45].

### v3

@@
 Our code is released at https://github.com/thu-uav/NeuralIMC.
+Index Terms:
+Robust Control, Reinforcement Learning, Sensorimotor Learning
 I Introduction
@@
 To separate the $\mathcal{L}_{1}$ estimator’s effect, we also include a privileged variant, DATT(GT), which directly uses ground-truth force disturbances in place of the estimator.
-As shown in Tab.
-II, our method consistently outperforms others across nearly all tasks, particularly under disturbances and out-of-distribution system dynamics.
-Compared to state-of-the-art learning-based methods, classical adaptive control methods perform worse due to their reliance on handcrafted parameters, which fail to generalize across diverse conditions.
-Among the baselines, DATT shows strong resilience to external disturbances, supported by its $\mathcal{L}_{1}$ estimator.
-However, its performance declines noticeably with unexpected dynamics in evaluation.
-Comparing DATT(GT) and DATT($\mathcal{L}_{1}$) reveals that the learned RL policy generalizes well when provided with ground-truth inputs, suggesting that the performance drop is primarily due to the handcrafted $\mathcal{L}_{1}$ estimator, which struggles to generalize across different systems.
-On the other hand.
-the teacher-student training approach in RMA results in the student policy underperforming compared to the teacher policy and shows instability under unseen systems, highlighting its unreliability for learning-based prediction.
-In contrast, our method employs the proposed explicit model to establish a baseline prediction of the expected body transformation, effectively detecting discrepancies caused by internal and/or external sources.
-This informative feedback enhances the control policy’s responsiveness.
-Consequently, our method outperforms all baselines, including privileged variants like DATT(GT) and RMA(T), likely due to the informative predictive error feedback.
+As shown in Tab. II, our method consistently outperforms others across nearly all tasks, particularly under disturbances and out-of-distribution (OOD) system dynamics.
+Among estimator-based baselines—both explicit (e.g., $\mathcal{L}_{1}$ estimator in $\mathcal{L}_{1}$-PID, $\mathcal{L}_{1}$-MPPI, DATT) and implicit (e.g., RMA’s neural encoder)—DATT is the most resilient, benefiting from its explicit $\mathcal{L}_{1}$ estimator coupled with an RL-based controller.
+Nevertheless, all estimator-based methods degrade sharply under our extreme out-of-distribution scenarios—where internal dynamics shift or sudden external forces occur—because their estimators become unreliable and control robustness is compromised.
+Estimator-free approaches such as CAL and PPO also fail to manage these varied disturbances.
+In contrast, our method computes the predictive error between expected and observed body transformations, directly signaling state drift and disturbance magnitude.
+A policy trained on this error maintains strong performance across both nominal and unexpected conditions.
+We also acknowledge that limited hyperparameter sweeps in our experiments may contribute to the baselines’ degraded performance.
+To ensure a fair comparison, we include privileged variants (DATT (GT) and RMA (T)) that receive ground-truth states rather than estimates, representing an upper tuning bound.
+Even against these privileged methods, our predictive error-driven controller achieves superior and more stable control, supporting our hypothesis that explicit predictive error feedback offers a simple, generalizable path to robust performance.
 TABLE III: Simulation results on quadrupedal robots, averaged over five random seeds.